# Solar photospheric velocities measured in space: a comparison between SO/PHI-HRT and SDO/HMI

### v2

@@ Acknowledgements.
We thank the anonymous referee for their comments and suggestions that improved the quality of this manuscript. Solar Orbiter is a space mission of international collaboration between ESA and NASA, operated by ESA. We are grateful to the ESA SOC and MOC teams for their support. The German contribution to SO/PHI is funded by the BMWi through DLR and by MPG central funds. The Spanish contribution is funded by AEI/MCIN/10.13039/501100011033/ and European Union “NextGenerationEU/PRTR” (RTI2018-096886-C5, PID2021-125325OB-C5, PCI2022-135009-2, PCI2022-135029-2) and ERDF “A way of making Europe”; “Center of Excellence Severo Ochoa” awards to IAA-CSIC (SEV-2017-0709, CEX2021-001131-S). The French contribution is funded by CNES. SDO is the first mission to be launched for NASA’s Living With a Star (LWS) Program. The HMI data are courtesy of NASA/SDO and the HMI science team. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101097844 — project WINSUN). This research used version 6.0.6 (10.5281/zenodo.15690707) of the SunPy open source software package (The SunPy Community et al. 2020).References
- Albert et al. (2023) Albert, K., Krivova, N. A., Hirzberger, J., et al. 2023, A&A, 678, A163
- Bai et al. (2024) Bai, X., Deng, Y., Zhang, H., et al. 2024, Sol. Phys., 299, 157
- Bailén et al. (2024) Bailén, F. J., Orozco Suárez, D., Blanco Rodríguez, J., et al. 2024, A&A, 681, A58
- Calchetti et al. (2023) Calchetti, D., Stangalini, M., Jafarzadeh, S., et al. 2023, A&A, 674, A109
- Castellanos Durán et al. (2021) Castellanos Durán, J. S., Lagg, A., & Solanki, S. K. 2021, A&A, 651, L1
- Couvidat et al. (2016) Couvidat, S., Schou, J., Hoeksema, J. T., et al. 2016, Sol. Phys., 291, 1887
- de la Cruz Rodríguez et al. (2011) de la Cruz Rodríguez, J., Kiselman, D., & Carlsson, M. 2011, A&A, 528, A113
- Evershed (1909) Evershed, J. 1909, MNRAS, 69, 454
- Finley et al. (2025) Finley, A. J., Brun, A. S., Strugarek, A., & Perri, B. 2025, A&A, 697, A217
- Gan et al. (2023) Gan, W., Zhu, C., Deng, Y., et al. 2023, Sol. Phys., 298, 68
- Gandorfer et al. (2018) Gandorfer, A., Grauf, B., Staub, J., et al. 2018, in Space telescopes and instrumentation 2018: Optical, infrared, and millimeter wave, Vol. 10698, SPIE, 1403–1415
- Grigor’ev et al. (2011) Grigor’ev, V. M., Ermakova, L. V., & Khlystova, A. I. 2011, Astronomy Reports, 55, 163
- Hathaway & Rightmire (2011) Hathaway, D. H. & Rightmire, L. 2011, ApJ, 729, 80
- Jafarzadeh et al. (2025) Jafarzadeh, S., Jess, D. B., Stangalini, M., et al. 2025, Nature Reviews Methods Primers, 5, 21
- Jess et al. (2023) Jess, D. B., Jafarzadeh, S., Keys, P. H., et al. 2023, Living Reviews in Solar Physics, 20, 1
@@
- Kahil et al. (2022) Kahil, F., Gandorfer, A., Hirzberger, J., et al. 2022, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 12180, Space Telescopes and Instrumentation 2022: Optical, Infrared, and Millimeter Wave, ed. L. E. Coyle, S. Matsuura, & M. D. Perrin, 121803F
- Kazachenko et al. (2015) Kazachenko, M. D., Fisher, G. H., Welsch, B. T., Liu, Y., & Sun, X. 2015, ApJ, 811, 16
- Kosugi et al. (2007) Kosugi, T., Matsuzaki, K., Sakao, T., et al. 2007, Sol. Phys., 243, 3
- Lites et al. (2013) Lites, B. W., Akin, D. L., Card, G., et al. 2013, Sol. Phys., 283, 579
- Lites et al. (1998) Lites, B. W., Skumanich, A., & Martinez Pillet, V. 1998, A&A, 333, 1053
- Loeschl et al. (2024a) Loeschl, P., Hirzberger, J., Solanki, S. K., Schou, J., & Valori, G. 2024a, A&A, 682, A108
- Loeschl et al. (2024b) Loeschl, P., Valori, G., Hirzberger, J., et al. 2024b, A&A, 681, A59
- Löhner-Böttcher & Schlichenmaier (2013) Löhner-Böttcher, J. & Schlichenmaier, R. 2013, A&A, 551, A105
- Löhner-Böttcher et al. (2019) Löhner-Böttcher, J., Schmidt, W., Schlichenmaier, R., Steinmetz, T., & Holzwarth, R. 2019, A&A, 624, A57
- Martínez Oliveros et al. (2011) Martínez Oliveros, J. C., Couvidat, S., Schou, J., et al. 2011, Sol. Phys., 269, 269
- Martínez Pillet et al. (1997) Martínez Pillet, V., Lites, B. W., & Skumanich, A. 1997, ApJ, 474, 810
- Mathew et al. (2003) Mathew, S. K., Lagg, A., Solanki, S. K., et al. 2003, A&A, 410, 695
- Moreno Vacas et al. (2024) Moreno Vacas, A., Orozco Suárez, D., Strecker, H., et al. 2024, A&A, 685, A28
- Müller et al. (2020) Müller, D., St. Cyr, O. C., Zouganelis, I., et al. 2020, A&A, 642, A1
- Oba et al. (2020) Oba, T., Iida, Y., & Shimizu, T. 2020, ApJ, 890, 141
- Okamoto et al. (2009) Okamoto, T. J., Tsuneta, S., Lites, B. W., et al. 2009, ApJ, 697, 913
- Orozco Suárez & Del Toro Iniesta (2007) Orozco Suárez, D. & Del Toro Iniesta, J. C. 2007, A&A, 462, 1137
- Pesnell et al. (2012) Pesnell, W. D., Thompson, B. J., & Chamberlin, P. C. 2012, Sol. Phys., 275, 3
- Romero Avila et al. (2024) Romero Avila, A., Inhester, B., Hirzberger, J., & Solanki, S. K. 2024, Sol. Phys., 299, 41
- Schlichenmaier et al. (2004) Schlichenmaier, R., Bellot Rubio, L. R., & Tritschler, A. 2004, A&A, 415, 731
- Schou et al. (2023) Schou, J., Hirzberger, J., Orozco Suárez, D., et al. 2023, A&A, 673, A84
- Schou et al. (2012) Schou, J., Scherrer, P. H., Bush, R. I., et al. 2012, Sol. Phys., 275, 229
- Schuck et al. (2016) Schuck, P. W., Antiochos, S. K., Leka, K. D., & Barnes, G. 2016, ApJ, 823, 101
- Sinjan et al. (2023) Sinjan, J., Calchetti, D., Hirzberger, J., et al. 2023, A&A, 673, A31
@@
- Sinjan et al. (2022) Sinjan, J., Calchetti, D., Hirzberger, J., et al. 2022, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 12189, Software and Cyberinfrastructure for Astronomy VII, 121891J
- Solanki et al. (2020) Solanki, S. K., del Toro Iniesta, J. C., Woch, J., et al. 2020, A&A, 642, A11
- Stief et al. (2019) Stief, F., Löhner-Böttcher, J., Schmidt, W., Steinmetz, T., & Holzwarth, R. 2019, A&A, 622, A34
- The SunPy Community et al. (2020) The SunPy Community, Barnes, W. T., Bobra, M. G., et al. 2020, The Astrophysical Journal, 890, 68
- Thompson (2006) Thompson, W. T. 2006, A&A, 449, 791
- Tsuneta et al. (2008) Tsuneta, S., Ichimoto, K., Katsukawa, Y., et al. 2008, Sol. Phys., 249, 167
- Valori et al. (2023) Valori, G., Calchetti, D., Moreno Vacas, A., et al. 2023, A&A, 677, A25
- Vernazza et al. (1981) Vernazza, J. E., Avrett, E. H., & Loeser, R. 1981, ApJS, 45, 635
- Vesa et al. (2023) Vesa, O., Jackiewicz, J., & Reardon, K. 2023, ApJ, 952, 58
- Vigeesh et al. (2017) Vigeesh, G., Jackiewicz, J., & Steiner, O. 2017, ApJ, 835, 148
- Welsch et al. (2013) Welsch, B. T., Fisher, G. H., & Sun, X. 2013, ApJ, 765, 98
- Zouganelis et al. (2020) Zouganelis, I., De Groof, A., Walsh, A. P., et al. 2020, A&A, 642, A3
@@ Appendix A Large scale velocity components
The treatment of the large scale velocity components (LSC) in the solar photosphere and their contribution to the LoS velocity field has been extensively treated in many papers (e.g. Schuck et al. 2016; Castellanos Durán et al. 2021). Information about the solar coordinate system can be found in Thompson (2006). Here, for completeness, we summarize the equations used in our analysis. The solar differential rotation is treated according to Hathaway & Rightmire (2011):
@@
The meridional flow is also treated according to Hathaway & Rightmire (2011):
@@
The variation from center to limb of the convective blueshift signal is adopted from Castellanos Durán et al. (2021). Despite the dependence of this effect on the spectral resolution (see Stief et al. 2019; Löhner-Böttcher et al. 2019), we are not able to fit the center to limb profile of this effect because SO/PHI-HRT is not a full-disk imager and we do not have continuous observations of the Sun (e.g. as in Löhner-Böttcher & Schlichenmaier 2013). This term is then modeled by Castellanos Durán et al. (2021) as:

### v3

@@ Acknowledgements.
We thank the anonymous referee for their comments and suggestions that improved the quality of this manuscript. Solar Orbiter is a space mission of international collaboration between ESA and NASA, operated by ESA. We are grateful to the ESA SOC and MOC teams for their support. The German contribution to SO/PHI is funded by the BMWi through DLR and by MPG central funds. The Spanish contribution is funded by AEI/MCIN/10.13039/501100011033/ and European Union “NextGenerationEU/PRTR” (RTI2018-096886-C5, PID2021-125325OB-C5, PCI2022-135009-2, PCI2022-135029-2) and ERDF “A way of making Europe”; “Center of Excellence Severo Ochoa” awards to IAA-CSIC (SEV-2017-0709, CEX2021-001131-S). The French contribution is funded by CNES. SDO is the first mission to be launched for NASA’s Living With a Star (LWS) Program. The HMI data are courtesy of NASA/SDO and the HMI science team. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101097844 — project WINSUN). This research used version 6.0.6 (10.5281/zenodo.15690707) of the SunPy open source software package (sunpy_community2020).References
- Intensity contrast of solar network and faculae close to the solar limb, observed from two vantage points. A&A 678, pp. A163. External Links: Document, 2309.01474, ADS entry Cited by: §1.
- Calibration and Performance of the Full-Disk Vector MagnetoGraph (FMG) on Board the Advanced Space-Based Solar Observatory (ASO-S). Sol. Phys. 299 (11), pp. 157. External Links: Document, ADS entry Cited by: §1.
- Determination of the SO/PHI-HRT wavefront degradation using multiple defocused images. A&A 681, pp. A58. External Links: Document, ADS entry Cited by: §2.
- Spectropolarimetric investigation of magnetohydrodynamic wave modes in the photosphere: First results from PHI on board Solar Orbiter. A&A 674, pp. A109. External Links: Document, 2304.14155, ADS entry Cited by: §2.
- How rare are counter Evershed flows?. A&A 651, pp. L1. External Links: Document, 2106.05592, ADS entry Cited by: Appendix A, Appendix A, §3.
- Observables Processing for the Helioseismic and Magnetic Imager Instrument on the Solar Dynamics Observatory. Sol. Phys. 291 (7), pp. 1887–1938. External Links: Document, 1606.02368, ADS entry Cited by: §2.
- Solar velocity references from 3D HD photospheric models. A&A 528, pp. A113. External Links: Document, 1101.2671, ADS entry Cited by: §4.2.
- Radial movement in sun-spots. MNRAS 69, pp. 454. External Links: Document, ADS entry Cited by: §4.3.
- A prolific solar flare factory: Nearly continuous monitoring of an active region nest with Solar Orbiter. A&A 697, pp. A217. External Links: Document, 2504.06345, ADS entry Cited by: §1.
- The Advanced Space-Based Solar Observatory (ASO-S). Sol. Phys. 298 (5), pp. 68. External Links: Document, ADS entry Cited by: §1.
- The high resolution telescope (hrt) of the polarimetric and helioseismic imager (phi) onboard solar orbiter. In Space telescopes and instrumentation 2018: Optical, infrared, and millimeter wave, Vol. 10698, pp. 1403–1415. Cited by: §1.
- The dynamics of photospheric line-of-sight velocities in emerging active regions. Astronomy Reports 55 (2), pp. 163–173. External Links: Document, ADS entry Cited by: §4.2, §5.
- Variations in the Axisymmetric Transport of Magnetic Elements on the Sun: 1996-2010. ApJ 729 (2), pp. 80. External Links: Document, 1010.1242, ADS entry Cited by: Appendix A, Appendix A.
- Wave analysis tools. Nature Reviews Methods Primers 5, pp. 21. External Links: Document, ADS entry Cited by: §4.1.
- Waves in the lower solar atmosphere: the dawn of next-generation solar telescopes. Living Reviews in Solar Physics 20 (1), pp. 1. External Links: Document, 2212.09788, ADS entry Cited by: §1, §6.
@@
- Image quality of data products of the high resolution telescope of the polarimetric and helioseismic imager. In Space Telescopes and Instrumentation 2022: Optical, Infrared, and Millimeter Wave, L. E. Coyle, S. Matsuura, and M. D. Perrin (Eds.), Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 12180, pp. 121803F. External Links: Document, ADS entry Cited by: §2.
- Photospheric Electric Fields and Energy Fluxes in the Eruptive Active Region NOAA 11158. ApJ 811 (1), pp. 16. External Links: Document, 1505.05974, ADS entry Cited by: §1.
- The Hinode (Solar-B) Mission: An Overview. Sol. Phys. 243 (1), pp. 3–17. External Links: Document, ADS entry Cited by: §1.
- The Hinode Spectro-Polarimeter. Sol. Phys. 283 (2), pp. 579–599. External Links: Document, ADS entry Cited by: §1.
- Vector magnetic fields of emerging solar flux. I. Properties at the site of emergence. A&A 333, pp. 1053–1068. External Links: ADS entry Cited by: §1.
- Synoptic maps from two viewpoints. Preparing for maps from SDO/HMI and SO/PHI data. A&A 682, pp. A108. External Links: Document, ADS entry Cited by: §1.
- A first rapid synoptic magnetic field map using SDO/HMI and SO/PHI data. A&A 681, pp. A59. External Links: Document, ADS entry Cited by: §1.
- Correlations between sunspots and their moat flows. A&A 551, pp. A105. External Links: Document, 1301.2434, ADS entry Cited by: Appendix A.
- Convective blueshifts in the solar atmosphere. III. High-accuracy observations of spectral lines in the visible. A&A 624, pp. A57. External Links: Document, 1901.07606, ADS entry Cited by: Appendix A.
- Imaging Spectroscopy of a White-Light Solar Flare. Sol. Phys. 269 (2), pp. 269–281. External Links: Document, 1012.0344, ADS entry Cited by: §2.
- Active Region Magnetic Fields. I. Plage Fields. ApJ 474 (2), pp. 810–842. External Links: Document, ADS entry Cited by: §5.
- Three dimensional structure of a regular sunspot from the inversion of IR Stokes profiles. A&A 410, pp. 695–710. External Links: Document, ADS entry Cited by: §4.2.
- Comparison of magnetic data products from Solar Orbiter SO/PHI-FDT and SDO/HMI. A&A 685, pp. A28. External Links: Document, ADS entry Cited by: §1.
- The Solar Orbiter mission. Science overview. A&A 642, pp. A1. External Links: Document, 2009.00861, ADS entry Cited by: §1, §1.
- Average Radial Structures of Gas Convection in the Solar Granulation. ApJ 890 (2), pp. 141. External Links: Document, 2001.03575, ADS entry Cited by: §5, §6.
- Prominence Formation Associated with an Emerging Helical Flux Rope. ApJ 697 (1), pp. 913–922. External Links: Document, 0904.0007, ADS entry Cited by: §4.2.
- The usefulness of analytic response functions. A&A 462 (3), pp. 1137–1145. External Links: Document, 1211.1502, ADS entry Cited by: §2.
- The Solar Dynamics Observatory (SDO). Sol. Phys. 275 (1-2), pp. 3–15. External Links: Document, ADS entry Cited by: §1.
- Photospheric Stereoscopy: Direct Estimation of Solar Surface-Height Variations. Sol. Phys. 299 (4), pp. 41. External Links: Document, ADS entry Cited by: §1.
- Two-dimensional spectroscopy of a sunspot. II. Penumbral line asymmetries. A&A 415, pp. 731–737. External Links: Document, ADS entry Cited by: §4.3.
- The ratio of horizontal to vertical displacement in solar oscillations estimated from combined SO/PHI and SDO/HMI observations. A&A 673, pp. A84. External Links: Document, 2303.16844, ADS entry Cited by: §4.1, §5, §5.
- Design and Ground Calibration of the Helioseismic and Magnetic Imager (HMI) Instrument on the Solar Dynamics Observatory (SDO). Sol. Phys. 275 (1-2), pp. 229–259. External Links: Document, ADS entry Cited by: §1.
- Achieving Consistent Doppler Measurements from SDO/HMI Vector Field Inversions. ApJ 823 (2), pp. 101. External Links: Document, 1511.06500, ADS entry Cited by: Appendix A, Appendix A, §3, §5.
- Magnetic fields inferred by Solar Orbiter: A comparison between SO/PHI-HRT and SDO/HMI. A&A 673, pp. A31. External Links: Document, 2303.16771, ADS entry Cited by: §1, §2.
@@
- The on-ground data reduction and calibration pipeline for SO/PHI-HRT. In Software and Cyberinfrastructure for Astronomy VII, Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 12189, pp. 121891J. External Links: Document, 2208.14904, ADS entry Cited by: §2.
- The Polarimetric and Helioseismic Imager on Solar Orbiter. A&A 642, pp. A11. External Links: Document, 1903.11061, ADS entry Cited by: §1.
- Convective blueshifts in the solar atmosphere. II. High-accuracy observations of the Fe I 6173.3 Å line and deviations of full-disk Dopplergrams. A&A 622, pp. A34. External Links: Document, 1811.08685, ADS entry Cited by: Appendix A, §5.
- Coordinate systems for solar image data. A&A 449 (2), pp. 791–803. External Links: Document, ADS entry Cited by: Appendix A.
- The Solar Optical Telescope for the Hinode Mission: An Overview. Sol. Phys. 249 (2), pp. 167–196. External Links: Document, 0711.1715, ADS entry Cited by: §1.
- Stereoscopic disambiguation of vector magnetograms: First applications to SO/PHI-HRT data. A&A 677, pp. A25. External Links: Document, 2307.09907, ADS entry Cited by: §1.
- Structure of the solar chromosphere. III. Models of the EUV brightness components of the quiet sun.. ApJS 45, pp. 635–725. External Links: Document, ADS entry Cited by: §5.
- Multiheight Observations of Atmospheric Gravity Waves at Solar Disk Center. ApJ 952 (1), pp. 58. External Links: Document, 2307.02426, ADS entry Cited by: §4.1.
- Internal Gravity Waves in the Magnetized Solar Atmosphere. I. Magnetic Field Effects. ApJ 835 (2), pp. 148. External Links: Document, 1612.04729, ADS entry Cited by: §4.1.
- A Magnetic Calibration of Photospheric Doppler Velocities. ApJ 765 (2), pp. 98. External Links: Document, 1201.2451, ADS entry Cited by: §4.2, §5.
- The Solar Orbiter Science Activity Plan. Translating solar and heliospheric physics questions into action. A&A 642, pp. A3. External Links: Document, 2009.10772, ADS entry Cited by: §2.
@@ Appendix A Large scale velocity components
The treatment of the large scale velocity components (LSC) in the solar photosphere and their contribution to the LoS velocity field has been extensively treated in many papers (e.g. Schuck et al. 2016; Castellanos Durán et al. 2021). Information about the solar coordinate system can be found in Thompson (2006). Here, for completeness, we summarize the equations used in our analysis. The solar differential rotation is treated according to Hathaway and Rightmire (2011):
@@
The meridional flow is also treated according to Hathaway and Rightmire (2011):
@@
The variation from center to limb of the convective blueshift signal is adopted from Castellanos Durán et al. (2021). Despite the dependence of this effect on the spectral resolution (see Stief et al. 2019; Löhner-Böttcher et al. 2019), we are not able to fit the center to limb profile of this effect because SO/PHI-HRT is not a full-disk imager and we do not have continuous observations of the Sun (e.g. as in Löhner-Böttcher and Schlichenmaier 2013). This term is then modeled by Castellanos Durán et al. (2021) as: